Haroldo, Xerxes, Regina, Shaindel, Murray, and Georgia are invited to a dinner party. They arrive in a random order and all arrive at different times. What is the probability that Xeres arrives first AND Regina arrives last?
Answer #1

Total number of possible ways = 6! = 720

Number of ways with Xeres arrives first AND Regina arrives last = 4! = 24

Hence,

P(Xeres arrives first AND Regina arrives last) = 24/720 = 1/30